# Implementation of the Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model in the GENIE neutrino event generator

### v1

@@ V Summary and Conclusion
The Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model has been implemented in the GENIE neutrino event generator. Even if this model allows a unified description of several channels (quasielastic, multinucleon emission and the coherent and incoherent single pion production) for both charged current and neutral current scattering, the present work concerns only the quasielastic and the multinucleon emission in the case of charged current (anti)neutrino scattering.
@@
Model predictions in GENIE for 12C, 16O and 40Ar have been compared with some available experimental measurements from T2K and MicroBooNE. Overall, a reasonable agreement has been observed.
This implementation of the Martini-Ericson-Chanfray-Marteau model in the widely used GENIE event generator is an important step towards eventual reduction of systematic uncertainties due to neutrino-nucleus cross section in a future combined analyses of data from the ongoing and next generation precision neutrino experiments.
@@ Acknowledgments
MM thanks Magda Ericson, Guy Chanfray, Jacques Marteau and Jean Delorme for all the work that led to the development of the model implemented in the present manuscript. Magda Ericson and Guy Chanfray are also acknowledged by MM for the long and continuous collaboration. The authors would also like to acknowledge support and feedback on this work from the T2K and GENIE collaborations.
@@ Appendix A Model predictions for 12C, 16O and 40Ca
In this Appendix we show the model predictions for the and charged-current total cross section as a function of the neutrino energy for three different nuclei which are of interest for the ongoing and future experiments: 12C (already presented and discussed in the main text), 16O and 40Ca (as a proxy for 40Ar). The corresponding comparisons are presented in Figs. 10 and 10 for and , respectively. A perfect agreement between the GENIE results and the original calculations is observed.

### v2

@@ V Summary and Conclusion
The Martini-Ericson-Chanfray-Marteau RPA-based (anti)neutrino cross-section model has been implemented in the GENIE neutrino event generator. Even if this model allows a unified description of several channels (quasielastic, multinucleon emission and the coherent and incoherent single pion production) for both charged- and neutral-current scattering, the present work concerns only the quasielastic and the multinucleon emission in the case of charged-current (anti)neutrino scattering. The implementation of single pion production as well as of electromagnetic and neutral-current scattering is left for future works.
@@
Model predictions in GENIE for 12C, 16O, and 40Ar have been compared with some available experimental measurements from T2K and MicroBooNE. Overall, a reasonable agreement has been observed.
This implementation of the Martini-Ericson-Chanfray-Marteau model in the widely used GENIE event generator is an important step toward eventual reduction of systematic uncertainties due to neutrino-nucleus cross section in future combined analyses of data from the ongoing and next generation precision neutrino experiments.
@@ Acknowledgments
M. M. thanks Magda Ericson, Guy Chanfray, Jacques Marteau and Jean Delorme for all the work that led to the development of the model implemented in the present manuscript. Magda Ericson and Guy Chanfray are also acknowledged by M. M. for the long and continuous collaboration. The authors would also like to acknowledge support and feedback on this work from the T2K and GENIE Collaborations.
@@ Appendix A Model predictions for 12C, 16O and 40Ca
In this appendix we show the model predictions for the and charged-current total cross section as a function of the neutrino energy for three different nuclei that are of interest for the ongoing and future experiments: 12C (already presented and discussed in the main text), 16O, and 40Ca (as a proxy for 40Ar). The corresponding comparisons are presented in Figs. 10 and 10 for and , respectively.
A perfect agreement between the GENIE results and the original calculations is observed.
Appendix B A brief summary of the npnh treatment in the Martini-Ericson-Chanfray-Marteau model
In this appendix we describe how npnh excitations are treated in the Martini et al. model by following an historical perspective which summarizes Sec. IV C of Ref. Martini et al. (2009) and Sec. II of Ref. Martini et al. (2010).
There are several sources of npnh excitations. One of these sources arises from the modification of the width in the nuclear medium. This effect was introduced and parametrized by Oset and Salcedo Oset and Salcedo (1987), which considered the two-body (2p2h) [Fig. 11(d)] and the three-body (3p3h) channels in the case of real pion or photon absorption. Martini et al. have used their parametrization of the modified width, although the kinematics of neutrino interaction is different since we are in the spacelike region, which could be a source of uncertainty. For the other terms not reducible to a modification of a width, of which some examples are shown in Fig. 11, they have first used, as in the papers of Marteau et al. Marteau (1999); Marteau et al. (2000), a parametrization of Delorme and Guichon Delorme and Guichon (1989) (called “old” in Sec. IV C of Ref. Martini et al. (2009)). They exploited a calculation by Shimizu and Faessler Shimizu and Faessler (1980) of the absorptive part of the p-wave pion-nucleus optical potential at threshold. It is known that the absorption mechanism of pions is a two-nucleon one, which means that in the final state two nucleons are ejected. This is a 2p2h excitation. In Ref. Shimizu and Faessler (1980), the absorption is described by three types of terms (see Fig. 11).
The first one [Fig. 11(a) with pion lines replacing the weak current ones as in the next diagrams] arises from the correlations, essentially from the tensor correlations. Another one involves a excitation [Fig. 11(c)]. The third one is an interference between the nucleon correlation and terms ( interference) as in Fig. 11(b). The pion absorption calculation of Shimizu and Faessler Shimizu and Faessler (1980) is performed for threshold pion, i.e., for a vanishing three-momentum and an energy , which does not correspond to the neutrino situation. Delorme and Guichon Delorme and Guichon (1989) have then introduced in each absorption graph the corresponding energy dependence to obtain the bare 2p2h responses to be inserted in the RPA chain needed for neutrino interaction. The explicit expressions of the corresponding propagators are given in Appendix B 1 of Ref. Martini et al. (2009). The shortcoming of these formulas is that the momentum dependence is completely ignored. For this reason in Ref. Martini et al. (2009) Martini et al. introduced also a different approach that exploits the microscopic calculation of Alberico et al. Alberico et al. (1984) specifically aimed at the evaluation of the 2p2h contribution to the isospin spin-transverse response, measured in inclusive scattering. Their basic graphs are similar to those of Shimizu and Faessler Shimizu and Faessler (1980). In principle, this way of evaluation is definitely more satisfactory since the kinematical variables are correctly incorporated, but the results of Alberico et al. Alberico et al. (1984) are available only for a limited set of energy and momenta. Martini et al. have extended this range to cover the neutrino one through an approximate extrapolation. We refer to Ref. Martini et al. (2009) for the details. Martini et al. hence obtained, starting from the calculations of Alberico et al., “new” and responses entering in the new and components of the hadron tensor. For the part, which is not well covered in Ref. Alberico et al. (1984), Martini et al. have kept the previous parametrization, which already presents a proper dependence owing to the contribution of the in-medium width Oset and Salcedo (1987). To summarize, in the 2p2h sector, Martini et al. include NN correlations, MEC contributions, and NN correlations-MEC interference. Concerning the MEC, they decided to consider only the -MEC and to discard the other contributions from the explicit calculation because these other contributions are peculiar to the external probe and they want a “universal” spin-isospin 2p2h response, to use in (and deduced from) different processes, like electron scattering and pion absorption. The -MEC contribution is the dominant one in MEC, as shown, for example, in Fig. 9 of De Pace et al. (2003).